# bibliography.bib

### v1

@@ V-B $\Lambda^{*}$ Optimization with Hill Climbing
The number of crossings of shortest paths between pairs in $\tilde{C}$ can serve as reasonable cost metric [molavi_dependency-aware_2024]. Consider Fig. 1c, where an optimal solution can be found by directly computing the shortest paths between the qubit vertices. Swapping the locations of two labels, e.g., 00 and $3$, however, would inevitably lead to a suboptimal solution because the respective shortest paths would cross. While for this small example the intuition reflects an exact solution, for larger instances counting the crossings of shortest paths will be a heuristic metric in general.
@@
Hill climbing [yang_nature-inspired_2014, gendreau_handbook_2019, jacobson_analyzing_2004], is a simple meta-heuristic that can find local extrema. The algorithm uses a random choice of a mapping $\Lambda^{*}$ as initial value and computes the metric for each so-called neighboring solution. The algorithm proceeds in a greedy fashion: Once the best solution among the neighboring ones is chosen based on the defined metric, the next neighborhood is searched. This iterative procedure is repeated until the number of maximal iterations is reached or no better solution in the neighborhood can be found. We use the simple adaption of random restarts, i.e., running the algorithm with multiple randomly chosen $\Lambda^{*}$ as initial values. This way a slightly larger search space can potentially be explored. Thus, the number of restarts and the maximal iterations per restart are important parameters of the optimization procedure.
The neighborhood for the greedy search in the hill climbing is defined as all combinations of swaps of the positions of each pair in $\tilde{C}$ [molavi_dependency-aware_2024]. Moreover, the factory positions are not changed during the hill climbing optimization. Merely the qubit labels, $\Lambda^{*}$ are optimized. Note, however, that one can potentially improve the method by employing more sophisticated algorithms such as ABC [hancer_artificial_2020] or genetic algorithms [lambora-genetic]. Our work is a proof-of-principle demonstration. We therefore leave a detailed comparison of different optimization strategies open for future work.
@@ VI Microscopic Substrate realizations
In this section, we outline the subsystem code formalism for lattice surgery introduced by Vuillot et al. [vuillot_code_2019] and use it to argue fault-tolerance properties of lattice surgery between distant logical patches on a substrate utilizing distance-preserving snakes. We also discuss microscopic realizations of the color code and surface code substrates.
@@ VI-B Color Code Substrate
For the color code substrate we choose a hexagonal tiling on which we define the regions $\mathcal{A}^{*}\cup\mathcal{L}\cup\mathcal{F}$ that form distance $d=2t+1$ triangular color code patches, where $t$ is a positive integer. For this substrate each patch has $q=3t^{2}+3t+1$ physical data qubits [bravyi_doubled_2015].
To perform lattice surgery between two neighboring logical patches, their adjacent boundaries are connected by measuring intermediary operators that create a semitransparent domain wall (STDW) [kesselring_anyon_2024]. For a $Z_{L}Z_{L}$ merge between two adjacent logical patches, one extends the existing $Z$ and $X$ stabilizers and uses further weight $2,3,5$, and $6$ stabilizers along the STDW. For a $X_{L}X_{L}$ merge, the roles of $X$ and $Z$ stabilizers must be swapped. For ease of presentation we focus on the case of measuring $Z_{L}Z_{L}$.
@@
The set of gauge operators $\mathcal{L}_{g}$ contains the new intermediary weight $3,5,$ and $6$ $Z$ plaquettes $\mathcal{I}$ along the STDWs next to logical patches, depicted as pink dotted plaquettes in  Fig. 6a. These are the gauge operators that will be fixed for the merge. The stabilizers in $S_{\text{split}}$ commute with the new plaquettes, except for the weight 4 $X$ checks along the boundaries of the logical patches, which are also in $\mathcal{L}_{g}$ (blue dotted plaquettes in Fig. 6a). These can be extended to weight 6 stabilizers of $\mathcal{Q}$ along the STDW. To ensure that the stabilizer group of $\mathcal{Q}$ has the correct form, we add $X$ string operators to $\mathcal{L}_{g}$ as anti-commuting conjugate partners of the operators in $\mathcal{I}$. To do this, it is helpful to distinguish an inner boundary of the patch, i.e., a boundary which is connected to another patch with a STDW and an outer boundary without adjacent connected patch. Starting from the left logical patch $L_{1}$, a string operator can be chosen on $L_{1}$ such that one of its two point-like excitations is on the outer boundary of $L_{1}$, and the other excitation appears on an element from $\mathcal{I}$ on an inner boundary. One can add another string operator extending through the snake that can create an excitation on an element of $\mathcal{I}$ on the inner boundary adjacent to the top right logical patch (cf. Fig. 6a). This choice of gauge operators ensures that they commute with the stabilizers of the subsystem code (grey faces) and anti-commute with elements of $\mathcal{I}$, which are not supposed to be stabilizers, as they anti-commute with the gauge operators of the split code (blue dotted faces).
To fix the gauge, the operators in $\mathcal{I}$ are measured and the stabilizers are updated according to the Gottesman-Knill theorem [gottesman1998heisenberg].
Note that the domain walls dictate a special behaviour of the logical string operators. Logical $Z_{L}$ strings can terminate at the STDW such that the weight of $Z_{L}$ operators remains the same as for the initial logical patches. $X$ strings cannot terminate at the STDW but only on the color code boundaries of the appropriate color [kesselring_anyon_2024]. Thus $\mathcal{Q}_{\text{merged}}$ and $\mathcal{Q}$ have logical $Z$ operators of weight $d$ of a single logical patch. Bare logical $X$ operators of $\mathcal{Q}$ can stretch from one logical patch to another, but since there are $X$ gauge operators supported on the left logical patch and the snake, the minimum distance of a dressed logical operator of $\mathcal{Q}$ is $d$. In the example in Fig. 6 the resulting operator is supported on the top right logical patch. Hence, the code distance of the subsystem code is at least the same as of an individual logical patch.
@@
Note that the color code lattice surgery scheme can also be done with $\mathcal{A}\backslash\mathcal{A}^{*}=\emptyset$, as proposed in [landahl_quantum_2014, kesselring_anyon_2024], but requires higher weight stabilizers.
@@ VI-C Surface Code Substrate
Figure 8: Valid paths on the surface code substrate with folded surface codes as logical data patches. Folded surface codes are displayed in light blue, snakes in grey, and logical ancilla patches in green. $\mathcal{R}$ is shown in orange and a routing path between logical qubits is marked in violet. (a) The shortest path on $\mathcal{R}$ between logical data patches 00 and $3$ is a valid path. (b) The shortest path on $\mathcal{R}$ between logical patch 1 and 2 is not valid, since the pair of triangles marked by red dashed lines would allow us to do either $X_{L}X_{L}$ or $Z_{L}Z_{L}$ lattice surgery, but not both as needed for the implementation of a $\mathrm{CNOT}$ gate. A slightly adapted valid path is displayed in violet together with a possible logical ancilla patch placement.
@@
We illustrate how to construct distance preserving snakes, by considering the surface code snake illustrated in Fig. 9, which connects a folded surface code patch and a logical surface code patch. Here, $S_{\text{split}}$ consists of the stabilizers of the disjoint logical patches and stabilizers on the ancilla space on the snake. $S_{\text{merged}}$ consists of one joint patch formed by additional checks in the rotated surface code part of the snake. The intermediary $Z$ checks $\mathcal{I}$ on the snake are supported on the weight 2 and 4 $Z$ checks that act on one of the qubits marked with orange dots in Fig. 9. These are in $\mathcal{L}_{g}$ and are again the ones we gauge fix for the merge. Moreover, the weight $3$ $X$ checks along the boundary of the logical ancilla patch and the folded surface code patch anti-commute with the new $Z$ checks formed on plaquettes in $\mathcal{I}$. These are also in $\mathcal{L}_{g}$ and can be extended to weight 4 faces that commute with the new $Z$ checks in $S$. Similarly to the color code substrate, one can choose $X$ string operators in $\mathcal{L}_{g}$ with support on the left logical patch and on the snake to be anti-commuting conjugate partners of elements in $\mathcal{I}$.
@@ VII Numerical Insights Into the Macroscopic Compilation With the Color Code
Here we describe numerical experiments to explore how the depth, number of available magic state patches, reset time of the magic state factories, and the chosen optimization metric influence the behavior of the proposed macroscopic compilation routine from Sec. V for the color code substrate.
@@ VII-A Setup
The macroscopic compilation operates on the logical routing graph $\mathcal{R}$. The valid paths on $\mathcal{R}$ for the color code substrate do not require special constraints, despite the fact that paths must be large enough to host at least one logical ancilla patch in order to perform a $\mathrm{CNOT}$ between logical data patches. This is ensured by the graph structure and the requirement that no direct connections between logical data patches are allowed. We consider three choices of layouts as illustrated in Fig. 5. Magic state patches are placed at the boundary of the routing graph and connected with two edges to the bulk of the routing graph, while the remaining edge is reserved for the realization of the magic state factory. We assume that for each factory a new $\mathrm{T}$ state is available after some reset time given as parameter.
Figure 10: Mean improvements from the hill climbing depending on different circuit types. Here, purely random $\mathrm{CNOT}{}$ circuits without T gates are considered. “seq.” means that at most two $\mathrm{CNOT}{}$s can be applied in parallel. The “rand.” circuits have $\mathrm{CNOT}{}$s drawn from a random distribution. “max.” means that all available qubits can be used in parallel, i.e., $q/2$ gates are applied in parallel. Per data point, 50 random circuits were sampled. Each run of the hill climbing algorithm uses 50 iterations for each of the 10 restarts. In total, $4q$ gates per circuit were used.
@@ VII-B $\mathrm{CNOT}{}$ Circuit Compilation
The comparison between different layouts shows that the optimized $\Lambda^{*}$ on the pair layout tends to be improved less, despite the “maximal” circuit type. A reason for this may be that sparser layouts with larger routing space encounter less crossings during the VDP solving such that one can assume less potential of $\Lambda^{*}$ improvement for them. However, if the initial parallelism is large enough—increasing the possibility of crossings—also optimization for the pair layout becomes fruitful. In contrast, for “maximal” and the row and hexagonal layouts, $\tilde{\Delta}$ is not considerably larger than for the “random” circuits. A reason for this may be that the heuristic metric cannot capture the system’s behaviour if too many crossings arise due to the high density for the layout together with high parallelism of the circuit. Too many “pushs” (cf. Sec. V) during the VDP solving cannot be reflected by the heuristic crossing metric.
Figure 11: Top row: Mean improvements $\tilde{\Delta}=(\Delta_{i}-\Delta_{f})/\Delta_{i}$. Bottom row: Absolute depth $\Delta_{f}$ after optimization. Left column: heuristic crossing metric. Right column: Exact routing metric. Per data point 50 different “random” circuits with $q=24$ qubits and $r=0.8$ were sampled. The “row” layout was chosen as well as the $\mathcal{R}$ for the color code substrate. Per circuit, $4q$ gates were considered.
@@ VII-C $\mathrm{CNOT}{}+\mathrm{T}$ Circuit Compilation
Here we consider $\mathrm{T}$ gates in the input circuits as well. To this end, random circuits are sampled in such a way that there will be a specific ratio of $\mathrm{CNOT}{}$ gates in comparison to the total gates $r=\#CNOT/(\#CNOT+\#T)$. This ratio is important as $\mathrm{T}$ gates make an efficient solution particularly difficult due to the reset time $t$ and the limited availability of magic state patches. We investigate the behavior of the hill climbing optimization with different numbers of magic state patches as well as different reset times. The results for “random” circuits are shown in Fig. 11. The first row displays the improvements $\tilde{\Delta}$ for the heuristic crossing metric (left) and the more expensive metric $\Delta_{f}$ of the shortest-first VDP solving (right). The improvements with the heuristic metric work best for many magic state patches and a small reset time, while it deteriorates for more challenging setups. Using the exact metric $\Delta_{f}$ makes the hill climbing solutions more advantageous, also for more challenging setups. The bottom row of Fig. 11 displays the corresponding absolute values for $\Delta_{f}$ of the VDP solving runs of the best $\Lambda^{*}$ from the hill climbing optimization. It is apparent that irrespective the chosen metric the achieved depth is the lowest for many magic state patches and low reset time $t$, with better results using the metric $\Delta_{f}$ for hill climbing.
@@ VIII Conclusion and Outlook
Our methods enable future optimizations of fault-tolerant compilation. At the microscopic level, we envision exploring different substrates, possibly motivated by different quantum computing platforms, and taking hardware imperfections into account. For instance, one possible direction is to incorporate into our framework the recently-developed methods for adapting surface code patches in the presence of defective qubits and gates [auger2017fault, debroy_luci_2024, leroux_snakes_2024]. At the macroscopic level, the $\Lambda^{*}$ optimization and algorithms to solve the (extended) VDP problem allow for further improvements; commuting some Clifford gates to the end of the circuit and exploring potential parallelism in the modified circuit may be valuable. Future work may also explore how to incorporate novel schemes for magic state distillation and cultivation [gidney_magic_2024], with a careful analysis of their layouts within the bulk of the routing graph.
@@ Acknowledgements
L.H., L.B., and R.W. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101001318). This work was part of the Munich Quantum Valley, which the Bavarian state government supports with funds from the Hightech Agenda Bayern Plus, and has been supported by the BMWK based on a decision by the German Bundestag through project QuaST, as well as by the BMK, BMDW, and the State of Upper Austria in the frame of the COMET program (managed by the FFG).
[…]

### v2

@@ V-B $\Lambda^{*}$ Optimization with Hill Climbing
The number of crossings of shortest paths between pairs in $\tilde{C}$ can serve as reasonable cost metric [13]. Consider Fig. 1c, where an optimal solution can be found by directly computing the shortest paths between the qubit vertices. Swapping the locations of two labels, e.g., 0 and 333, however, would inevitably lead to a suboptimal solution because the respective shortest paths would cross. While for this small example the intuition reflects an exact solution, for larger instances counting the crossings of shortest paths will be a heuristic metric in general.
@@
Hill climbing [55, 56, 57], is a simple meta-heuristic that can find local extrema. The algorithm uses a random choice of a mapping $\Lambda^{*}$ as initial value and computes the metric for each so-called neighboring solution. The algorithm proceeds in a greedy fashion: Once the best solution among the neighboring ones is chosen based on the defined metric, the next neighborhood is searched. This iterative procedure is repeated until the number of maximal iterations is reached or no better solution in the neighborhood can be found. We use the simple adaption of random restarts, i.e., running the algorithm with multiple randomly chosen $\Lambda^{*}$ as initial values. This way a slightly larger search space can potentially be explored. Thus, the number of restarts and the maximal iterations per restart are important parameters of the optimization procedure.
The neighborhood for the greedy search in the hill climbing is defined as all combinations of swaps of the positions of each pair in $\tilde{C}$ [13]. Moreover, the factory positions are not changed during the hill climbing optimization. Merely the qubit labels, $\Lambda^{*}$ are optimized. Note, however, that one can potentially improve the method by employing more sophisticated algorithms such as ABC [58] or genetic algorithms [59]. Our work is a proof-of-principle demonstration. We therefore leave a detailed comparison of different optimization strategies open for future work.
@@ VI Microscopic Substrate realizations
In this section, we outline the subsystem code formalism for lattice surgery introduced by Vuillot et al. [60] and use it to argue fault-tolerance properties of lattice surgery between distant logical patches on a substrate utilizing distance-preserving snakes. We also discuss microscopic realizations of the color code and surface code substrates.
@@ VI-B Color Code Substrate
For the color code substrate we choose a hexagonal tiling on which we define the regions $\mathcal{A}^{*}\cup\mathcal{L}\cup\mathcal{F}$ that form distance $d=2t+1$ triangular color code patches, where $t$ is a positive integer. For this substrate each patch has $q=3t^{2}+3t+1$ physical data qubits.
To perform lattice surgery between two neighboring logical patches, their adjacent boundaries are connected by measuring intermediary operators that create a semitransparent domain wall (STDW) [61]. For a $Z_{L}Z_{L}$ merge between two adjacent logical patches, one extends the existing $Z$ and $X$ stabilizers and uses further weight 2,3,52,3,52 , 3 , 5, and 666 stabilizers along the STDW. For a $X_{L}X_{L}$ merge, the roles of $X$ and $Z$ stabilizers must be swapped. For ease of presentation we focus on the case of measuring $Z_{L}Z_{L}$.
@@
The set of gauge operators $\mathcal{L}_{g}$ contains the new intermediary weight 3,5,3,5,3 , 5 , and 666 $Z$ plaquettes $\mathcal{I}$ along the STDWs next to logical patches, depicted as pink dotted plaquettes in  Fig. 6a. These are the gauge operators that will be fixed for the merge. The stabilizers in $S_{\text{split}}$ commute with the new plaquettes, except for the weight 4 $X$ checks along the boundaries of the logical patches, which are also in $\mathcal{L}_{g}$ (blue dotted plaquettes in Fig. 6a). These can be extended to weight 6 stabilizers of $\mathcal{Q}$ along the STDW. To ensure that the stabilizer group of $\mathcal{Q}$ has the correct form, we add $X$ string operators to $\mathcal{L}_{g}$ as anti-commuting conjugate partners of the operators in $\mathcal{I}$. To do this, it is helpful to distinguish an inner boundary of the patch, i.e., a boundary which is connected to another patch with a STDW and an outer boundary without adjacent connected patch. Starting from the left logical patch $L_{1}$, a string operator can be chosen on $L_{1}$ such that one of its two point-like excitations is on the outer boundary of $L_{1}$, and the other excitation appears on an element from $\mathcal{I}$ on an inner boundary. One can add another string operator extending through the snake that can create an excitation on an element of $\mathcal{I}$ on the inner boundary adjacent to the top right logical patch (cf. Fig. 6a). This choice of gauge operators ensures that they commute with the stabilizers of the subsystem code (grey faces) and anti-commute with elements of $\mathcal{I}$, which are not supposed to be stabilizers, as they anti-commute with the gauge operators of the split code (blue dotted faces).
To fix the gauge, the operators in $\mathcal{I}$ are measured and the stabilizers are updated according to the Gottesman-Knill theorem [28].
Note that the domain walls dictate a special behaviour of the logical string operators. Logical $Z_{L}$ strings can terminate at the STDW such that the weight of $Z_{L}$ operators remains the same as for the initial logical patches. $X$ strings cannot terminate at the STDW but only on the color code boundaries of the appropriate color [61]. Thus $\mathcal{Q}_{\text{merged}}$ and $\mathcal{Q}$ have logical $Z$ operators of weight $d$ of a single logical patch. Bare logical $X$ operators of $\mathcal{Q}$ can stretch from one logical patch to another, but since there are $X$ gauge operators supported on the left logical patch and the snake, the minimum distance of a dressed logical operator of $\mathcal{Q}$ is $d$. In the example in Fig. 6 the resulting operator is supported on the top right logical patch. Hence, the code distance of the subsystem code is at least the same as of an individual logical patch.
@@
Note that the color code lattice surgery scheme can also be done with $\mathcal{A}\backslash\mathcal{A}^{*}=\emptyset$, as proposed in [18, 61], but requires higher weight stabilizers.
@@ VI-C Surface Code Substrate
Figure 8: Valid paths on the surface code substrate with folded surface codes as logical data patches. Folded surface codes are displayed in light blue, snakes in grey, and logical ancilla patches in green. $\mathcal{R}$ is shown in orange and a routing path between logical qubits is marked in violet. (a) The shortest path on $\mathcal{R}$ between logical data patches 0 and 333 is a valid path. (b) The shortest path on $\mathcal{R}$ between logical patch 1 and 2 is not valid, since the pair of triangles marked by red dashed lines would allow us to do either $X_{L}X_{L}$ or $Z_{L}Z_{L}$ lattice surgery, but not both as needed for the implementation of a $\mathrm{CNOT}$ gate. A slightly adapted valid path is displayed in violet together with a possible logical ancilla patch placement.
@@
We illustrate how to construct distance preserving snakes, by considering the surface code snake illustrated in Fig. 9, which connects a folded surface code patch and a logical surface code patch. Here, $S_{\text{split}}$ consists of the stabilizers of the disjoint logical patches and stabilizers on the ancilla space on the snake. $S_{\text{merged}}$ consists of one joint patch formed by additional checks in the rotated surface code part of the snake. The intermediary $Z$ checks $\mathcal{I}$ on the snake are supported on the weight 2 and 4 $Z$ checks that act on one of the qubits marked with orange dots in Fig. 9. These are in $\mathcal{L}_{g}$ and are again the ones we gauge fix for the merge. Moreover, the weight 333 $X$ checks along the boundary of the logical ancilla patch and the folded surface code patch anti-commute with the new $Z$ checks formed on plaquettes in $\mathcal{I}$. These are also in $\mathcal{L}_{g}$ and can be extended to weight 4 faces that commute with the new $Z$ checks in $S$. Similarly to the color code substrate, one can choose $X$ string operators in $\mathcal{L}_{g}$ with support on the left logical patch and on the snake to be anti-commuting conjugate partners of elements in $\mathcal{I}$.
@@ VII Numerical Insights Into the Macroscopic Compilation With the Color Code
Here we describe numerical experiments to explore how the depth, number of available magic state patches, reset time of the magic state factories, and the chosen optimization metric influence the behavior of the proposed macroscopic compilation routine from Sec. V for the color code substrate. The implementation is part of the Munich Quantum Toolkit [62].
@@ VII-A Setup
The macroscopic compilation operates on the logical routing graph $\mathcal{R}$. The valid paths on $\mathcal{R}$ for the color code substrate do not require special constraints, despite the fact that paths must be large enough to host at least one logical ancilla patch in order to perform a $\mathrm{CNOT}$ between logical data patches. This is ensured by the graph structure and the requirement that no direct connections between logical data patches are allowed. We consider three choices of layouts as illustrated in Fig. 5. Magic state patches are placed at the boundary of the routing graph and connected with two edges to the bulk of the routing graph, while the remaining edge is reserved for the realization of the magic state factory. We assume that for each factory a new $\mathrm{T}$ state is available after some time interval given as parameter.
Figure 10: Mean improvements from the hill climbing depending on different circuit types. Here, purely random $\mathrm{CNOT}{}$ circuits without T gates are considered. “seq.” means that at most two $\mathrm{CNOT}{}$s can be applied in parallel. The “rand.” circuits have $\mathrm{CNOT}{}$s drawn from a random distribution. “max.” means that all available qubits can be used in parallel, i.e., $q/2$ gates are applied in parallel. Per data point, 50 random circuits with $q=24$ qubits were sampled. Each run of the hill climbing algorithm uses 50 iterations for each of the 10 restarts. In total, $4q$ gates per circuit were used.
@@ VII-B $\mathrm{CNOT}{}$ Circuit Compilation
The comparison between different layouts shows that the optimized $\Lambda^{*}$ on the pair layout tends to be improved less. A reason for this may be that sparser layouts with larger routing space encounter less crossings during the VDP solving such that one can assume less potential of $\Lambda^{*}$ improvement for them. However, if the initial parallelism is large enough—increasing the possibility of crossings—also optimization for the pair layout becomes more fruitful. Overall, $\tilde{\Delta}$ increases with growing parallelism for all layouts. The asymptotic packing ratio of the row and hexagonal layout are very close which is why they show similar behavior.
[…]
@@ VII-C $\mathrm{CNOT}{}+\mathrm{T}$ Circuit Compilation
Here we consider $\mathrm{T}$ gates in the input circuits as well. To this end, random circuits are sampled in such a way that there will be a specific ratio of $\mathrm{CNOT}{}$ gates in comparison to the total gates $r=\#CNOT/(\#CNOT+\#T)$. This ratio is important as $\mathrm{T}$ gates make an efficient solution particularly difficult due to the reset time $t$ and the limited availability of magic state patches. We investigate the behavior of the hill climbing optimization with different numbers of magic state patches as well as different reset times. The results for “random” circuits are shown in Fig. 11. The first row displays the improvements $\tilde{\Delta}$ for the heuristic crossing metric (a) and the more expensive metric $\Delta_{f}$ of the shortest-first VDP solving (b). The improvements work best for many magic state patches and a small reset time, while they are most likely reduced for more challenging setups. Using the exact metric $\Delta_{f}$ makes the hill climbing solutions more advantageous than with the heuristic. The bottom row of Fig. 11 displays the corresponding absolute values for $\Delta_{f}$ of the VDP solving runs of the best $\Lambda^{*}$ from the hill climbing optimization for the respective metrics (c, d). It is apparent that irrespective the chosen metric the achieved depth is the lowest for many magic state patches and low reset time $t$, with slightly better results using the metric $\Delta_{f}$ for hill climbing.
@@ VIII Conclusion and Outlook
Our methods enable future optimizations of fault-tolerant compilation. At the microscopic level, we envision exploring different substrates, possibly motivated by different quantum computing platforms, and taking hardware imperfections into account. For instance, one possible direction is to incorporate into our framework the recently-developed methods for adapting surface code patches in the presence of defective qubits and gates [63, 64, 65]. At the macroscopic level, the $\Lambda^{*}$ optimization and algorithms to solve the (extended) VDP problem allow for further improvements; commuting some Clifford gates to the end of the circuit and exploring potential parallelism in the modified circuit may be valuable. Future work may also explore how to incorporate novel schemes for magic state distillation and cultivation [66], with a careful analysis of their layouts within the bulk of the routing graph.
@@ Acknowledgements
L.H., L.B., and R.W. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101001318). This work was part of the Munich Quantum Valley, which the Bavarian state government supports with funds from the Hightech Agenda Bayern Plus as well as by the BMK, BMDW, and the State of Upper Austria in the frame of the COMET program (managed by the FFG).